Determine the number of protons, neutrons, and ele
Answer #1

a)
atomic number of Cr = 24
so,
number of proton = 24
number of elctron = 24
number of neutron = 52-24 =28

b)
atomic number of Cd = 48
so,
number of proton = 48
number of elctron = 48
number of neutron = 112-48 =64

c)
atomic number of Ba = 56
so,
number of proton = 56
number of elctron = 56 - 2 = 54
number of neutron = 137-56 = 81

d)
atomic number of Cu= 29
so,
number of proton = 29
number of elctron = 28
number of neutron = 63-29 =34

e)
atomic number of fe = 26
so,
number of proton = 26
number of elctron = 26-3= 23
number of neutron = 55-26 =29